A time-varying net force acting on a 4.0-kg particle causes the particle to have a


displacement given by x = 2.0 t - 3.0 t2 + 1.0 t3


, where x is in meters and t is in seconds.


(a) Use Work- energy theorem to find the work done on the particle in the first 3.0 s of motion.


(b) The average power delivered in the first 3 s.


(c) the instantaneous power delivered at t = 2 s.

1
Expert's answer
2021-12-10T12:22:47-0500

(a) "v(t)=2-6t+3t^2" ; "t=0\to v_0=2\ (m/s)"

"t=3\ s\to v=11\ (m/s)"


"W=\Delta KE=mv^2/2-mv_0^2/2=(m/2)(v^2-v_0^2)="


"=(4/2)(11^2-2^2)=234\ (J)"


(b) "P_{avr}=W/t=234/3=78\ (W)"


(c) "P(2)=F\cdot v=m(6t-6)(2-6t+3t^2)="


"=4\cdot(6\cdot2-6)(2-6\cdot2+3\cdot2^2)=48\ (W)"